# Why is Grain Growth Not Curvature Flow?

### v1

@@ Abstract
Grain growth describes the increase in the mean grain size with time during the annealing of a polycrystal; it is widely accepted that this is driven by capillarity (surface tension). Although classically modeled and interpreted as mean curvature flow, a growing body of evidence has demonstrated that this is overly simplistic and inconsistent with experimental observations. Despite many proposals for the source of this failure, it remains a subject of ongoing debate. By considering the fundamental microscopic mechanism of grain boundary (GB) migration, we demonstrate that internal stress (shear coupling) associated with GB motion is the primary factor causing deviations in grain growth from mean curvature flow. This is achieved through the application of a two-dimensional bicrystallography-respecting continuum model for GB migration to grain growth, implemented in a diffuse-interface approach. These results are validated by comparison with atomistic simulations and the scatter in the curvature-velocity relationship is shown to be in excellent agreement with experimental results for grain growth in 3D polycrystals (not possible based on other proposed explanations). This work also demonstrates how to account for the mechanical effects that are integral to grain growth in crystalline solids.
I Introduction
@@
Polycrystalline materials consist of numerous single crystals (grains) with different crystallographic orientations separated by interfaces/grain boundaries (GBs). Polycrystals are ubiquitous in both natural and man-made material; most engineering materials are polycrystalline. The microstructure, namely the geometry of the grain ensemble and GB network, plays a major role in establishing the (physical, mechanical,…) properties of polycrystalline materials. The classical theory of microstructure evolution focuses on the relaxation of the excess energy of a GB network, by GB migration resulting in grain annihilation. GBs tend to migrate in the direction that reduces their area, that is towards their center of curvature. This gives rise to the well-known description of grain growth as mean curvature ($\kappa$) flow; the GB velocity may be expressed as $v=M\gamma\kappa$, where $M$ is the GB mobility and $\gamma$ is the (isotropic) GB energy [1, 2, 3, 4]. Although this model describes the evolution of similar-looking microstructures, such as relatively dry soap froths, it neglects the fundamental nature of the material in the grains (crystals rather than gases). The crystalline nature of the grains implies that the properties are anisotropic and, importantly, the material itself supports mechanical deformation, i.e., stress and strain.
@@
Several recent investigations clearly show that a mean curvature flow description of grain growth fails to explain experimental and atomistic simulation observations. Phenomena that point to this failure include the existence of stress-driven grain growth [5, 6, 7], grain rotation [8, 9, 10], GB sliding [11, 12], and peculiar grain-size distributions [13, 14]. Recent experimental measurements and molecular dynamics (MD) simulations found that, in some systems, GB velocity is not correlated with GB mean curvature [15, 16, 17]. Moreover, many observations show that GB migration is accompanied by shear across the GB both in bicrystals and polycrystals; i.e., shear-coupled migration associated with defect (disconnection) motion along GBs [18, 19, 20].
@@
Figure 1: 2D grain growth simulations. (a) Initial microstructure and orientation $\theta$ field of a typical 1000 grain polycrystal. (b) Mean grain size $R^{2}$ versus time $t$ during grain growth. Each solid and thick curve is averaged over 30 independent PF simulations (i.e., the light and thin curves) with and without GB shear coupling. Microstructure evolution (c) without GB shear coupling (mean curvature flow) and (d) with shear coupling (shear modulus is that of Al).
@@
Shear-coupled migration plays a key role in accounting for much of the phenomenology associated with GB motion. The coupling factor, $\beta=v_{\parallel}/v_{\perp}$, relates the GB migration rate $v_{\perp}$ and the GB shear rates $v_{\parallel}$. When shear-coupled GB migration is constrained (e.g., by triple junctions), an internal stress field is generated [20, 21]. The stress, in turn, affects GB migration and influences grain growth [21]. We recently unified these elastic and curvature effects in a bicrystallography-respecting continuum GB migration model [22, 23] and demonstrated that GB shear coupling naturally leads to internal stress generation [23, 24], GB faceting transition [25], and grain rotation [26] during GB migration in bicrystal geometries. Nonetheless, two critical questions remain: (1) how does shear-coupling influence large-scale microstructure evolution? and (2) can shear-coupling account for the recent experimental observations that grain growth is not curvature flow? Fortunately, both simulation and experimental observations have reached a level where detailed comparisons may now be made at the microscopic and statistical levels to quantitively answer these questions.
@@
In this work, we demonstrate that the internal stresses generated by shear-coupled GB motion is the main cause for the poor correlations between the GB velocity and the curvature during grain growth in polycrystals. We do this through a series of large-scale grain growth simulations. These simulations are performed using the bicrystallography-respecting continuum model for GB migration mentioned above, implemented in a diffuse-interface approach [23, 24] extending the classical phase field (PF) model of grain growth [27, 28]. Simulated microstructure evolution, accounting for both GB curvature effect and internal stress, reproduces the scatter in the curvature-velocity relationship reported in several recent experimental and atomistic simulation studies. We also demonstrate that other proposed explanations for the failure of the application of curvature flow to grain growth are insufficient to account for the recent experimental observations.
II Results
II.1 Effect of internal stress on 2D grain growth
@@
We first discuss a series of PF simulations on grain growth of 2D polycrystals performed with and without GB shear coupling to identify how shear coupling modifies essential features of microstructure evolution; modeling details are in the Methods section and Supplementary Information (SI). The initial polycrystalline microstructures for the grain growth simulations are based on PF simulations that are nearly equivalent to Voronoi tessellations of a Poisson point process with 1000 grains (see [24]). We assume that the 2D polycrystals are composed of [110]-textured face-centered cubic grains (there are two sets of closely-packed/low-energy GB planes, perpendicular to each other, in the CSL lattices). If we choose the reference system such that $[001]\,||\,\mathbf{e}_{x}$, $[1\bar{1}2]\,||\,\mathbf{e}_{y}$ and $[110]\,||\,\mathbf{e}_{z}$, crystal symmetry dictates that all grain orientation angles are $\theta\in$[0,$\pi$); see the orientation field of a typical initial microstructure in Fig. 1a. For simplicity, all GB energies are assumed to be the same. Therefore, evolution without shear coupling reduces to classical mean curvature flow. Next, we consider the effect of GB shear coupling by assigning shear coupling factors ($\beta$) to each GB according to the GB misorientation angle [29, 20] (see SI for details on how shear coupling factors are assigned).
@@
Figure 2: Internal stress fields generated by GB migration. (a) A schematic of a grain in a polycrystal with GB shear coupling (shear modulus is roughly that of Al) caused by disconnection motion on each GB; see the zoomed schematic in (b). The evolving internal stress field $\sigma_{\text{12}}$ during grain growth at a grain size of (c) $1\times 10^{5}$ and (d) $2\times 10^{5}$ see Fig. 1(d) and (e) its distribution at different times. (f)-(h) Same as (c)-(e) for the von Mises stress.
@@
Figure 1b shows the evolution of the mean grain (area) size ($R^{2}$) with and without GB shear coupling. Each solid, thick curve is the averaged result of 30 independent PF simulations (i.e., the light, thin curves) with different initial microstructures (i.e., different random initial microstructures and orientation distributions). These results show that the parabolic grain growth law is valid, with or without shear coupling. The grain growth rate is, however, $\sim 50\%$ larger with shear coupling than without, pointing to the importance of elastic effects. Figures 1c,d show the evolution of the microstructure starting from the same initial configuration, Fig. 1a. Shear coupling leads to significant differences in the temporal evolution of the microstructures; i.e., both the GB network and grain orientations.
@@
Shear-coupled GB migration is mediated by the flow of disconnections–i.e., line defects with both step and dislocation character constrained to and lying on GBs [30, 31, 20]. Thus, GB shear coupling can lead to the generation of internal stress. Figure 2a shows a schematic of disconnection flow and the resultant shear along the GBs delineating a single grain in a polycrystal. Focussing on the GB segment inside the dashed rectangle in Fig. 2b, we see that macroscopically, curved GBs can be described in terms of microscopic/atomistic disconnections. The step character is represented by the horizontal and vertical black lines and the dislocation character of the disconnections are indicated in green. Assuming that this GB segment is migrating downward (i.e., toward its mean curvature/capillarity), such GB migration occurs through the motion of the two vertical steps toward each other (see the green arrows, $\mathbf{v}_{\rm d}$). At the same time, the two oppositely signed dislocations (Burgers vectors) glide towards each other, leading to a relative shear across the GB; this is the origin/mechanism of shear coupling. As Burgers vector flow along each GB, a spatially and temporally varying internal stress field is produced.
@@
Figures 2c,d show the internal stress $\sigma_{12}$ distribution for the microstructures in Fig. 1d. The generated internal stress constantly evolves and is spatially inhomogeneous within the whole microstructure and each grain. Figure 2e shows the evolution of the fractions of $\sigma_{12}$; the lightest and darkest lines are for the initial and final microstructures, respectively. The $\sigma_{12}$ stress distributions are nearly Gaussian with widths that decrease during grain growth. To eliminate the effect of the coordinate system and investigate the elastic energy density evolution, we also show the von Mises stresses in Figs. 2f,g for the same microstructures. This stress distribution is also inhomogeneous and evolving. The von Mises stresses are nearly Rayleigh distributed, with peaks moving to lower stress during microstructure evolution (Fig. 2h). Both Figs. 2e and 2h demonstrate that internal stresses are relieved by grain growth; i.e., grain growth dissipates elastic energy.
@@
Note that stresses observed in the simulations shown in Fig. 2 exhibit maximum stresses as large as $\sim\mu/10$, which are comparable to the theoretical strength of the material. Where these large stresses are highly localized near the triple junctions (TJs), their origin is related to the constrained GB sliding at TJs. We also note that there can be very large stresses within grain interiors as well. In most metals, such large shear stress could be relieved by some plastic deformation within the grains. While the possibility of bulk plastic deformation is not included in these simulations, these results suggest that some microplasticity (dislocation, twin) may result from GB migration in real polycrystals [32, 33, 17].
@@
II.2 Failure of the von Neumann-Mullins relation
@@
where $R^{2}$ is the grain area, $M$ and $\gamma$ are the isotropic GB mobility and GB energy, $N$ is the number of neighbors of the grain and $\alpha_{i}$ is the dihedral angle at a triple junction; in the isotropic case all $\alpha_{i}=\pi/3$. This implies that grains with more/less than six neighbors will grow/shrink.
Figure 3a shows grain area change vs. the number of neighbors of individual grains obtained in our 2D PF grain growth simulations (shown in Fig. 1) in the absence of GB shear coupling. The thick blue line is the ideal grain growth relation (Eq. 1), and unsurprisingly, all data points are clustered very close to this line. Almost all grains with less than 6 neighbors shrink, while those with more than 6 grow, consistent with the von Neumann-Mullins relation (6 neighbors implies that the mean GB curvature is zero). While the von Neumann-Mullins grain growth relation is exact, the small discrepancy in Fig. 3a is associated with the diffuse boundary nature of the simulations. However, when the simulations include GB shear coupling, very large deviations from the ideal grain growth relation are observed; c.f., Figs. 3a and 3b. When the same simulation are performed with shear coupling, we note that the grain growth rate data has a very large standard deviation. This implies that the number of neighbors is no longer a good predictor of whether a grain will shrink or grow; in other words, the von Neumann-Mullins relationship fails in the presence of shear coupling. Nonetheless, on average, grains with a large number of neighbors tend to grow and few neighbors tend to shrink.
In 3D, curvature flow simulations (no shear coupling) [35] show that the grain volume growth rate versus number of neighbors is similar to the 2D curvature flow simulation results (i.e., the standard deviation of the data is small and there is a critical $N$ above which grains tend to grow and below which they shrink); c.f., Figs. 3a and 3c. However, grain growth experiments on 3D SrTiO3 show that the standard deviation of the grain volume growth rate is very large; see Fig. 3d. The 3D curvature flow simulation results and the 3D experiments differ because the experiments (unavoidably) include shear coupling while the 3D curvature flow simulations do not. This is exactly analogous to the 2D results where simulations with and without shear coupling differ in the same manner. [^1] Note, the extension of the von Neumann-Mullins relation to 3D shows that the grain size growth rate should be proportional to the mean width of the grain [4] not the number of neighbors; unfortunately, such data is not available from the experiments [35].
@@
Figure 3: Failure of the von Neumann-Mullins relation. Scatter plots of individual grain area changes $\delta R^{2}/\delta t$ versus the number of neighbors $N$ during grain growth (a) without shear coupling (i.e., mean curvature flow) and (b) with shear coupling from 2D PF simulations. Reported grain volume change $\delta R^{3}/\delta t$ vs. the number of neighbors in 3D from (c) curvature flow grain growth simulations [35] and (d) a polycrystalline SrTiO3 grain growth experiment [35]. (c) and (d) are reproduced with permission from (Copyright 2023 Elsevier) The gray shaded regions in the plots indicate the 95% confidence intervals from the data shown.
II.3 Failure of mean curvature flow

### v3

# Why Grain Growth is Not Curvature Flow

@@ Abstract
Grain growth in polycrystals is traditionally considered a capillarity-driven process, where grain boundaries (GBs) migrate toward their centers of curvature (i.e., mean curvature flow) with a velocity proportional to the local curvature (including extensions to account for anisotropic GB energy and mobility). Experimental and simulation evidence shows that this simplistic view is untrue. We demonstrate that the failure of the classical mean curvature flow description of grain growth mainly originates from the shear deformation naturally coupled with GB motion (i.e., shear coupling). Our findings are built on large-scale microstructure evolution simulations incorporating the fundamental (crystallography-respecting) microscopic mechanism of GB migration. The nature of the deviations from curvature flow revealed in our simulations is consistent with observations in recent experimental studies on different materials. This work also demonstrates how to incorporate the mechanical effects that are essential to the accurate prediction of microstructure evolution.
@@
Polycrystalline materials consist of numerous single crystals (grains) with different crystallographic orientations separated by interfaces/grain boundaries (GBs). Polycrystals are ubiquitous in both natural and man-made materials; most engineering materials are polycrystalline. The microstructure, namely the geometry of the grain ensemble and GB network, plays a major role in establishing the (physical, mechanical,…) properties of polycrystalline materials. The classical theory of microstructure evolution focuses on the relaxation of the excess energy of a GB network by GB migration, resulting in grain annihilation. GBs tend to migrate in the direction that reduces their area, that is towards their center of curvature. This gives rise to the well-known description of grain growth as mean curvature ($\kappa$) flow; the GB velocity may be expressed as $v=M\gamma\kappa$, where $M$ is the GB mobility and $\gamma$ is the (isotropic) GB energy [1, 2, 3, 4]. Although this model describes the evolution of similar-looking microstructures, such as relatively dry soap froths, it neglects the fundamental nature of the material in the grains (crystals rather than gases). The crystalline nature of the grains implies that GB properties are anisotropic. This is often accounted for by extending the classical mean curvature flow description to the case of anisotropic GB mobility and/or energy depending on GB macroscopic degrees of freedom; i.e., weighted mean curvature flow [5, 6]. Importantly, the material itself also supports mechanical deformation, i.e., stress and strain.
@@
Several recent investigations clearly show that a (weighted) mean curvature flow description of grain growth fails to explain experimental and atomistic simulation observations. Phenomena that point to this failure include the existence of stress-driven grain growth [7, 8, 9], grain rotation [10, 11, 12], GB sliding [13, 14], grain growth acceleration under oscillatory fields [15], and peculiar grain-size distributions [16, 17]. Recent experimental measurements and molecular dynamics (MD) simulations found that, in some systems, GB velocity is not correlated with GB mean curvature [18, 19, 20]. Moreover, many observations show that GB migration is accompanied by shear across the GB, both in bicrystals and polycrystals; i.e., shear-coupled migration associated with defect (disconnection) motion along GBs [21, 22, 23, 15].
@@
Figure 1: 2D grain growth simulations. (a) Initial microstructure and orientation $\theta$ field of a typical 1000 grain polycrystal. (b) Mean grain size $\langle R^{2}\rangle$ versus time $t$ during grain growth. Each solid and thick curve is averaged over 30 independent PF simulations (i.e., the light and thin curves) with and without GB shear coupling. Microstructure evolution (c) without GB shear coupling (mean curvature flow) and (d) with shear coupling (shear modulus is that of Al).
@@
Shear-coupled migration plays a key role in accounting for much of the phenomenology associated with GB motion. The coupling factor, $\beta=v_{\parallel}/v_{\perp}$, relates the GB migration rate $v_{\perp}$ and the GB shear rate $v_{\parallel}$. When shear-coupled GB migration is constrained (e.g., by triple junctions), an internal stress field is generated [23, 24]. The stress, in turn, affects GB migration and influences grain growth [24]. We recently unified these elastic and curvature effects in a bicrystallography-respecting continuum GB migration model [25, 26] and demonstrated that GB shear coupling naturally leads to internal stress generation [26, 27], GB faceting transition [28], and grain rotation [29] during GB migration in bicrystal geometries. Nonetheless, two critical questions remain: (1) how does shear-coupling influence large-scale microstructure evolution? and (2) can shear-coupling account for the recent experimental observations that grain growth is not curvature flow? Fortunately, both simulation and experimental observations have reached a level where detailed comparisons may now be made at the microscopic and statistical levels to quantitively answer these questions.
@@
In this work, we demonstrate that the internal stresses generated by shear-coupled GB motion is the main cause for the poor correlations between the GB velocity and the curvature during grain growth in polycrystals. This is achieved through a series of large-scale grain growth simulations. They are performed using the bicrystallography-respecting continuum model for GB migration mentioned above, implemented in a diffuse-interface approach [26, 27] extending the classical phase field (PF) model of grain growth [30, 31]. Simulated microstructure evolution, accounting for both GB curvature effect and internal stress, reproduces the scatter in the curvature-velocity relationship reported in several recent experimental and atomistic simulation studies. We also demonstrate that other proposed explanations for the failure of the application of curvature flow to grain growth are insufficient to account for the recent experimental observations.
I Results
I.1 Effect of internal stress on 2D grain growth
@@
We first discuss a series of PF simulations on grain growth of 2D polycrystals performed with and without GB shear coupling to identify how shear coupling modifies essential features of microstructure evolution; modeling details are in the Methods section. The initial polycrystalline microstructures for the grain growth simulations are based on PF simulations that are nearly equivalent to Voronoi tessellations of a Poisson point process with 1000 grains (see [27]). We assume that the 2D polycrystals are composed of [110]-textured face-centered cubic grains (there are two sets of closely-packed/low-energy GB planes, perpendicular to each other, in the coincidence-site lattices). If we choose the reference system such that $[001]\,||\,\mathbf{e}_{x}$, $[1\bar{1}2]\,||\,\mathbf{e}_{y}$, and $[110]\,||\,\mathbf{e}_{z}$, crystal symmetry dictates that all grain orientation angles are $\theta\in$[0,$\pi$); see the orientation field of a typical initial microstructure in Fig. 1a. For simplicity, all GB energies are assumed to be the same. Therefore, evolution without shear coupling reduces to classical mean curvature flow. Next, we consider the effect of GB shear coupling by assigning shear coupling factors ($\beta$) to each GB according to the GB misorientation angle [32, 23]; see the Supporting Information (SI) for further details on how shear coupling factors are assigned.
@@
Figure 2: Internal stress fields generated by GB migration. (a) A schematic of a grain in a polycrystal with GB shear coupling caused by disconnection motion on each GB; see the zoomed schematic in (b). The evolving internal stress field $\sigma_{\text{12}}$ during grain growth at a grain size of (c) $1\times 10^{5}$ and (d) $2\times 10^{5}$ see Fig. 1(d) and (e) its distribution at different times (shear modulus is roughly that of Al). (f)-(h) Same as (c)-(e) for the von Mises stress.
@@
Figure 1b shows the evolution of the mean grain (area) size ($\langle R^{2}\rangle$) with and without GB shear coupling. Each solid, thick curve is the averaged result of 30 independent PF simulations (i.e., the light, thin curves) with different initial microstructures (i.e., different random initial microstructures and orientation distributions). These results show that the parabolic grain growth law is approximately valid with or without shear coupling. The grain growth rate is, however, $\sim 50\%$ larger with shear coupling than without, pointing to the importance of elastic effects. Figures 1c,d show the evolution of the microstructure starting from the same initial configuration, Fig. 1a. Shear coupling leads to significant differences in the temporal evolution of the microstructures: i.e., both the GB network and grain orientations.
@@
Shear-coupled GB migration is mediated by the flow of disconnections–i.e., line defects with both step and dislocation character constrained to and lying on GBs [33, 34, 23]. Thus, GB shear coupling can lead to the generation of internal stress. Figure 2a shows a schematic of disconnection flow and the resultant shear along the GBs delineating a single grain in a polycrystal. Focussing on the GB segment inside the dashed rectangle in Fig. 2b, we see that macroscopically curved GBs can be described in terms of microscopic/atomistic disconnections. The step character is represented by the horizontal and vertical black lines, and the dislocation character of the disconnections is indicated in green. Assuming that this GB segment is migrating downward (i.e., toward its mean curvature/capillarity), such GB migration occurs through the motion of the two vertical steps toward each other (see the green arrows). At the same time, the two oppositely signed dislocations (Burgers vectors) glide towards each other, leading to a relative shear across the GB; this is the origin/mechanism of shear coupling. As Burgers vector flow along each GB, a spatially and temporally varying internal stress field is produced.
@@
Figures 2c,d show the internal stress $\sigma_{12}$ distribution for the microstructures in Fig. 1d. The generated internal stress constantly evolves and is spatially inhomogeneous within the whole microstructure and each grain. Figure 2e shows the evolution of the fractions of $\sigma_{12}$; the lightest and darkest lines are for the initial and final microstructures, respectively. The $\sigma_{12}$ stress distributions are nearly Gaussian, with widths that decrease during grain growth. To eliminate the effect of the coordinate system and investigate the elastic energy density evolution, we also show the von Mises stresses in Figs. 2f,g for the same microstructures. This stress distribution is also inhomogeneous and evolving. The von Mises stresses are nearly Rayleigh distributed, with peaks moving to lower stress during microstructure evolution (Fig. 2h). Both Figs. 2e and 2h demonstrate that internal stresses are relieved by grain growth; i.e., grain growth dissipates elastic energy.
Figure 3: Failure of the von Neumann-Mullins relation. Scatter plots of individual grain area changes $\delta R^{2}/\delta t$ versus the number of neighbors $N$ during grain growth (a) without shear coupling (i.e., mean curvature flow) and (b) with shear coupling from 2D PF simulations. Reported grain volume change $\delta R^{3}/\delta t$ vs. the number of neighbors in 3D from (c) curvature flow grain growth simulations [35] and (d) a polycrystalline SrTiO3 grain growth experiment [35]. (c) and (d) are reproduced with permission from [35] (Copyright 2023 Elsevier) The gray shaded regions in the plots indicate the 99% confidence intervals from the data shown.
@@
Note that stresses observed in the simulations shown in Fig. 2 exhibit maximum stresses as large as $\sim\mu/10$, which are comparable to the theoretical strength of the material. Where these large stresses are highly localized near the triple junctions (TJs), the origin of these high stresses is related to the constrained GB sliding at TJs. We also note that there can be very large stresses within grain interiors as well. In most metals, such large shear stress could be relieved by some plastic deformation within the grains. While the possibility of bulk plastic deformation is not included in these simulations, these results suggest that some microplasticity (dislocation, twin) may result from GB migration in real polycrystals [36, 37, 20].
@@
I.2 Failure of the von Neumann-Mullins relation
@@
where $R^{2}$ is the grain area, $M$ and $\gamma$ are the isotropic GB mobility and GB energy, $N$ is the number of neighbors of the grain, $\alpha_{i}$ is the dihedral angle at a triple junction; in the isotropic case $\alpha_{i}=\pi/3$. This implies that grains with more/less than six neighbors will grow/shrink.
Figure 3a shows grain area change vs. the number of neighbors of individual grains obtained in our 2D PF grain growth simulations (shown in Fig. 1) in the absence of GB shear coupling. The thick blue line is the ideal grain growth relation (1), and unsurprisingly, all data points are clustered very close to this line. Almost all grains with less than six neighbors shrink, while those with more than six grow, consistent with the von Neumann-Mullins relation (having six neighbors implies that the mean GB curvature is zero). While the von Neumann-Mullins grain growth relation is exact, the small discrepancy in Fig. 3a is associated with the diffuse boundary nature of the simulations. However, when the simulations include GB shear coupling, very large deviations from the ideal grain growth relation are observed; $c.f.$, Figs. 3a and b. In this case, we note that the grain growth rate data has a very large standard deviation. This implies that the number of neighbors is no longer a good predictor of whether a grain will shrink or grow; in other words, the von Neumann-Mullins relationship fails in the presence of shear coupling. Nonetheless, on average, grains with a large number of neighbors tend to grow, and those with few neighbors tend to shrink.
In 3D, curvature flow simulations (no shear coupling) [35] show that the grain volume growth rate versus number of neighbors is similar to the 2D curvature flow simulation results (i.e., the standard deviation of the data is small and there is a critical $N$ above which grains tend to grow and below which they shrink); $c.f.$, Figs. 3 a and c. However, grain growth experiments on 3D SrTiO3 show that the standard deviation of the grain volume growth rate is very large; see Fig. 3d. The 3D curvature flow simulation results and the 3D experiments differ because the experiments (unavoidably) include shear coupling, while the 3D curvature flow simulations do not. This is exactly analogous to the 2D results, where simulations with and without shear coupling differ in the same manner. [^1] Note, the extension of the von Neumann-Mullins relation to 3D shows that the grain size growth rate should be proportional to the mean width of the grain [4], not the number of neighbors; unfortunately, such data is not available from the experiments [35].
@@
I.3 Failure of mean curvature flow